OpenStudy (anonymous):

If y = 4x-2 all over -5x then what is the value of x when y = -1 over 5??

OpenStudy (anonymous):

So just plug in -1/5 for y in your original problem

OpenStudy (anonymous):

y =(4x-2)/-5x -1/5=4x-2 1.8=4x .45=x

OpenStudy (anonymous):

thx for explaining it to me.. so what is the final answer in fraction form??

OpenStudy (dumbcow):

i disagree x = 2/3 you forgot about the -5x

OpenStudy (dumbcow):

-1/5 = (4x-2)/-5x multiply by -5x on both sides x = 4x-2 -3x = -2 x=2/3

OpenStudy (anonymous):

oh thank you dumbcow! I did forget about the -5 my bad !

OpenStudy (dumbcow):

no prob did you follow that sparky?

OpenStudy (anonymous):

yah I got it.. that makes sense now and that was one of the multiple choice answers... thx